OpenStudy (anonymous):

I need to prove that ln(1-x) = - sum of x^n/n from 1 to infinity, but I don't understand where the negative comes from. Help?

OpenStudy (anonymous):

Have you tried finding a pattern in the Taylor polynomial terms?

OpenStudy (anonymous):

No, because we haven't learned that yet.

OpenStudy (anonymous):

$$f(x)=\log(1-x)\\f'(x)=\frac{-1}{1-x}=-\frac1{1-x}\\f''(x)=-\frac{-1}{(1-x)^2}=-\frac1{(1-x)^2}\\f'''(x)=\frac{-2}{(1-x)^3}=-\frac2{(1-x)^3}\\f^{(4)}(x)=-\frac6{(1-x)^4}\\\dots$$At \(x=0\) observe these reduce to:$$f(0)=0\\f'(0)=-1\\f''(0)=-1\\f'''(0)=-2\\f^{(4)}(0)=-6\\\dots$$the key is to realize that the product rule of differentiation keeps multiplying the previous powers; e.g. for \(f^{(4)}\) our numerator \(6\) comes from \(3\times2\times1=3!\). In general, \(f^{(n)}(0)=(n-1)!\) by this pattern

OpenStudy (anonymous):

rather \(f^{(n)}(0)=-(n-1)!\) **

OpenStudy (anonymous):

From our Taylor expansion form at \(x=0\) it is obvious what happens next:$$f(x)=\sum_{n=0}^\infty\frac{f^{(n)}(0)}{n!}x^n=-\sum_{n=1}^\infty\frac{(n-1)!}{n!}x^n=-\sum_{n=1}^\infty\frac1nx^n$$

OpenStudy (anonymous):

notice we don't include the \(n=0\) term since \(f(0)=0\)

OpenStudy (anonymous):

Is there a way without using the Taylor expansion?

OpenStudy (anonymous):

I think I have an idea. Did you learn about the derivative/integral of a sum?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The problem itself is about a Taylor expansion :-p but @SithsAndGiggles has a brilliant idea

OpenStudy (anonymous):

tabby says he/she hasn't learned taylor expansions/series yet, though.

OpenStudy (anonymous):

Good effort nonetheless @oldrin.bataku

OpenStudy (anonymous):

Since \(f(x)=\ln(1-x)\), you have \(f'(x)=-\dfrac{1}{1-x}\). Note that for \(|x|<1\), \[-\frac{1}{1-x}=-\sum_{n=0}^\infty x^n\] Take the integral of the sum and you get \[f(x)=\ln(1-x)=\int\left(-\sum_{n=0}^\infty x^n\right)~dx\\ \ln(1-x)=\cdots\]

OpenStudy (anonymous):

Btw, I'm a girl.

OpenStudy (anonymous):

So what would the first few partial sums look like?

OpenStudy (anonymous):

Partial sums? Just how rigorous is this proof supposed to be? I would think the integral of the series would suffice to show that the given series represents \(f(x)\).